7x+5y=-6
4x+3y=-4

Answer by mananth(16946)

7 x + 5 y = -6 .............1
4 x + 3 y = -4 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 5
-21 x -15 y = 18
20 x 15 y = -20
Add the two equations
-1 x = -2
/ -1
x = 2
plug value of x in (1)
7 x + 5 y = -6
14 + 5 y = -6
5 y = -6 -14
5 y = -20
y = -4